Help please thank you.
Mai is flying a plane at an altitude of 1600 ft. She sights a stadium at an angle of depression of 35 degrees. What is Mai's approximate horizontal distance from the stadium?
A. 676 Feet? B. 1120 feet ? C. 1450 feet? D. 2285 feet?

Answer by Boreal(15235)

The line of sight is the hypotenuse of a right triangle. One leg is the vertical distance from the ground, the other leg is the horizontal distance, what is wanted, and the angle of 35 degrees has a tangent=1600/x
tan35=1600/x
x=1600/tan 35
tan 35 is 0.700
x=2285.03 feet
D is the answer
Draw a diagram (which I can't do here). It is very difficult to do this without a picture.